OpenStudy (anonymous):

Write the quantities 3x10^-4cm and 3x10^4km in ordinary notation.

OpenStudy (cwrw238):

10^-4 = 0.0001 x 3 = 0.0003 3 * 10^4 = 3*10*10*10*10 = 30,000 3 + 4 0's